OpenStudy (anonymous):

The graph of a quadratic function results in a: a. Hyperbola b. Parabola c. Ellipse d. Conical Ellipse

OpenStudy (anonymous):

@nirmalnema :o Do you know this

OpenStudy (nirmalnema):

parabola

OpenStudy (solomonzelman):

you asked this question yesterday.

OpenStudy (solomonzelman):

And I told you, that it is a parabola, that is opening up or down.

OpenStudy (solomonzelman):

Give me any quadratic equation, and I'll make it into a parabola.

OpenStudy (anonymous):

Why does it turn into a parabola

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

It turns into a U or ∩ shaped parabola, because the equation of the parabola is y=a(x-h)²+k (where (h,k) is the center. I would call it the minimum point of a U shaped, and the maximum point of a ∩ shaped parabola). Take any quadratic equation, rx²+sx+t=0 (where s and t are numbers) r(x²+{s/r}x)+t=0 r(x²+{s/r}x+N)-{N × r} +t=0 (where N is the number that would make parenthesis a perfect square. It is hard to explain it just like this without any example, but you propose an example, and I'll show.